OpenStudy (anonymous):

A car averages 27 miles per gallon. If gas costs $4.04 per gallon, which of the following is closest to how much the gas would cost for this car to travel 2,727 typical miles? A. $ 44.44 B. $109.08 C. $118.80 D. $408.04 E. $444.40 When x = 3 and y = 5, by how much does the value of 3x2 – 2y exceed the value of 2x2 – 3y ? F. 4 G. 14 H. 16 J. 20 K. 50 What is the value of x when 2x + 3 = 3x – 4 ? A. –7 B. negative one fifth C. 1 D. one fifth

OpenStudy (anonymous):

1) 2727/27=101 101*4.04=408.04 (B) 2) 3(3)^2-2(5)= 9^2-10= 71 and 2(3)^2-3(5)= 6^2-15=21 so: 71-21=50 (K) 3) 2x+3=3x-4 2x=3x-7 x=7 That wasn't an option you gave so I guess I would go with -7?